How many grams of hcl, hcl are required to prepare 4 L of 5M HCL in water?
1
Expert's answer
2019-05-23T07:58:21-0400

At first we will count the amount of HCl:

n = c * V = 5 M * 4 L = 20 mol;

Now we can find out the required mass of HCl:

m = Mr * n = 36.5 * 20 = 730 (g)